# Short-term CO2 emissions forecasting: insight from the Italian electricity market

### v1

@@ 3 Methodology
This section describes the setup of single models, forecast combinations, and the overall design of the forecast experiment. In the following, the hourly observations are denoted by $E_{1},\ldots,E_{t},\ldots,E_{n}$, where $t$ indexes the hours. To indicate observation at hour $h$ of day $i$, also the notation $E_{i,h}$ is used. The relationship between $t$, $i$, and $h$ is defined by $t=24(i-1)+h$, with $t=1,\ldots,8760$, $i=1,\ldots,365$, and $h=1,\ldots,24$.
@@ 3.1 Single model setup
In contrast, more complex and widely used time series models can capture the underlying temporal structures and seasonal patterns. A standard SARIMAX$(p,d,q)(P,D,Q)_{S}$ model is defined as follows:
$\phi(B)\Phi(B^{S})(1-B)^{d}(1-B^{S})^{D}E_{t}=\mu+\theta(B)\Theta(B^{S})+\alpha_{1}W_{t}+\alpha_{2}Y_{t}+\alpha_{3}B_{t}+\epsilon_{t}$ (3)
where: $B$ is the backshift operator, i.e. such that $B^{k}E_{t}=E_{t-k}$; $S$ denotes the seasonal period; $d$ and $D$ represent the orders of non-seasonal and seasonal differencing, respectively; $\phi$ is the vector of non-seasonal autoregressive coefficients; $\Phi_{k}$ is the vector of seasonal autoregressive coefficients; $\mu$ is a constant term; $\theta_{j}$ is the vector of non-seasonal moving average coefficients; and $\Theta_{l}$ is the vector of seasonal moving average coefficients; $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ represent the effects of the day of the week $(W_{t})$, intra-annual seasonality $(Y_{t})$, and bank holidays $(B_{t})$, respectively; $\epsilon_{t}$ is a white noise error term. SARIMAX models are identified by minimising their hourly average RMSE in the calibration set.
From a physical perspective, the processes involved in the production of carbon emissions from electricity generation are continuous in nature. Therefore, continuous-time stochastic processes offer a natural and potentially more accurate framework for modelling the evolution of carbon emissions. The adoption of functional models is motivated by the continuous nature of the underlying physical phenomenon (Ramsay and Silverman,, 2006). The functional approach enables the recovery of information on emissions released within the intervals between discrete hourly observations, capturing intra-hour dynamics that would otherwise be overlooked. Carbon emissions from electricity generation can be naturally represented as curves, making it logical to capture the serial dependence among daily curves and their temporal dynamics (Damon and Guillas,, 2002; Chen et al.,, 2021; Damon and Guillas,, 2024). Figure 3 shows the daily Italian carbon emission curves for the period 2021-2023. The raw data were transformed into functional data using the 12 Fourier basis (Ramsay and Silverman,, 2006).
Within this functional approach, let $\mathcal{H}$ be a real separable Hilbert space and consider the continuous-time carbon emissions process $E_{i}(\tau)$ defined in $\mathcal{H}$ over a time domain $\tau\in\mathrm{T}$, i.e. $E_{i}(\tau)\in L^{2}[1,24]$ with $t\in\mathbb{N}$ indicating the $i$-th daily profile. The seasonal functional autoregressive model with exogenous input SFARX($p$) is specified as follows
$E_{i}(\tau)=\sum_{i=1}^{p}\rho_{j}(E_{i-j}(\tau))+\gamma(E_{i-7}(\tau))+\alpha_{1}(W_{i}(\tau))+\alpha_{2}(Y_{i}(\tau))+\alpha_{3}(B_{i}(\tau))+\epsilon_{i}(\tau),$ (4)
where: $\rho_{j}(\cdot)$, $\gamma(\cdot),\alpha_{1}(\cdot),\alpha_{2}(\cdot)$ and $\alpha_{3}(\cdot)$ are bounded linear operators defined in $\mathcal{H}$. Specifically, the linear operator $\rho_{j}(\cdot)$ quantifies the influence of the lagged emission curve $E_{i-j}(\tau)$, while $\gamma(\cdot)$ quantifies the influence of the lagged emission curve of the previous week $E_{i-7}(\tau)$. The operators $\alpha_{1}(\cdot)$, $\alpha_{2}(\cdot)$, and $\alpha_{3}(\cdot)$ account for the effects of the weekly and yearly seasonalities and bank holidays, respectively. $W_{i}(\tau)$, $Y_{i}(\tau)$ and $B_{i}(\tau)$ are continuous exogenous variables defined in $\mathcal{H}$. The term $\epsilon_{i}(\tau)$ is a functional white noise process valued at $\mathcal{H}$. SFARX($p$) models are identified at both the aggregate level and for each zone based on their hourly average RMSE in the calibration set.
@@
Although the SFARX model accounts for the serial correlation between the functional emission curves by modelling it as a linear operator in the functional space (Bosq,, 2000), non-parametric approaches such as GAM are also used to capture potential non-linearities and provide greater flexibility in modelling relationships (Hastie and Tibshirani,, 1986; Hastie,, 2015). The GAM is defined by
$\begin{split}E_{t}=&f_{1}(T_{t},\lambda_{1})+f_{2}(Y_{t},\lambda_{2})+f_{3}(H_{t},\lambda_{3})+f_{4}(W_{t},\lambda_{4})+f_{5}(E_{t-168},\lambda_{5})\\ &+f_{6}(E_{t-24},\lambda_{6})+f_{7}(E_{t-48},\lambda_{7})+f_{8}(E_{t-72},\lambda_{8})+\alpha_{1}B_{t}+\epsilon_{t},\end{split}$ (5)
$f_{i}(\cdot,\lambda_{i})$ are smoothing spline functions with $\lambda_{i}$ degrees of freedom. In particular, $f_{1}(T_{t},\lambda_{1}=5)$ is a smoothing spline function with 5 degrees of freedom that captures the long-term trend in carbon emissions over time; $f_{2}(Y_{t},\lambda_{2}=5)$ is a periodic smoothing spline function with 5 degrees of freedom that models the monthly seasonal variation; $f_{3}(H_{t},\lambda_{3})$ accounts for intra-day periodicity; $f_{4}(W_{t},\lambda_{4})$ and $f_{5}(E_{t-168},\lambda_{5})$ capture the deterministic and stochastic components of the weekly periodicity, respectively; $f_{6}(E_{t-24},\lambda_{6})$, $f_{7}(E_{t-48},\lambda_{7})$, and $f_{8}(E_{t-72},\lambda_{8})$ account for autoregressive effects at 24, 48, and 72 hours (i.e., one to three days prior); $\alpha_{1}B_{t}$ represents the effect of bank holidays. To refer to model (5) and its more parsimonious specifications, the (non-standard) notation GAM($p$,$P$) is used, where $p$ and $P$ represent the number of daily and weekly lags included in the model, respectively. All other variables are always present. With this notation, model (5) is known as GAM(3,1). More parsimonious GAM specifications have been considered and selected according to the hourly average RMSE.
The TBATS model, introduced by Livera et al., (2011), is another popular and flexible time series forecasting method, designed to handle multiple seasonalities, trend components, and short-term dependencies (Tawiah et al.,, 2023; Stanislaus et al.,, 2024). The TBATS model (Hyndman and Khandakar,, 2008; Hyndman et al.,, 2025) can be mathematically represented as follows:
$E_{t}^{(\omega)}=l_{t-1}+\phi b_{t-1}+\sum_{i=1}^{M}S^{(i)}_{t-m_{i}}+d_{t},\text{ with }E_{t}^{(\omega)}=\begin{cases}\frac{E_{t}^{\omega}-1}{\omega},&\text{if }\omega\neq 0\\ \log(E_{t}),&\text{if }\omega=0\end{cases}$ (6)
where: $E_{t}^{(\omega)}$ is the Box-Cox transformation of $E_{t}$ with parameter $\omega$; $l_{t}=l_{t-1}+\phi b_{t-1}+\alpha d_{t}$ is the local level component; $b_{t}=(1-\phi)b+\phi b_{t-1}+\beta d_{t}$ is the short-term trend component; $\phi$ is the trend damping parameter; $S_{t}^{(i)}$ denotes the $i$th seasonal component at time $t$ modelled with a Fourier-based trigonometric representation; $d_{t}$ is the error term which follows an ARMA($p,q$) process; $m_{i}$ denotes the seasonal periods.
Alongside TBATS, a "semi-functional" approach such as K-nearest neighbours (KNN) is also considered for its flexibility in leveraging similarity across daily carbon emissions profiles (Lora et al.,, 2007; Tajmouati et al.,, 2024). Given the carbon emissions recorded up to day $i-1$, the goal is to predict the daily emission profile for day $i$. Let $\mathbf{E}_{i}=(E_{i,1},\ldots,E_{i,h},\ldots,E_{i,24})$ denote the vector of hourly carbon emissions for a generic day $i$, where $E_{i,h}$ is the emission recorded at hour $h$ of day $i$. Let $D(i,j)$ define the discrepancy between the daily emission profiles of days $i$ and $j$. $D(i,j)$ is evaluated using the Euclidean distance as follows:
@@
The KNN method identifies the $k$ nearest neighbours of day $i-1$, forming the neighbour set $NS=\{\mathbf{E}^{(1)},\mathbf{E}^{(2)},\ldots,\mathbf{E}^{(k)}\}$, where $\mathbf{E}^{(1)}$ is the nearest day in the past, $\mathbf{E}^{(2)}$ the second nearest, and $\mathbf{E}^{(k)}$ the k-th nearest. The 24-hour emission profile for day $i$ is then predicted as a weighted average of the emissions on the days following those of the neighbour set:
@@
where $w_{k}\in[0,1]$ is a weight inversely proportional to the distance between $\mathbf{E}^{(k)}$ and $\mathbf{E}_{t-1}$. The optimal $k$ is chosen by minimising the hourly average RMSE (Hnin et al.,, 2024). An example is provided by the elbow plot in Figure 4 for Italy.
@@ 3.2 Forecast combination setup
The idea of combining forecasts dates back to Bates and Granger, (1969), who demonstrated that forecast combinations typically outperform individual models (Clemen,, 1989; Surowiecki,, 2005; Timmermann,, 2006). Basically, different forecast combinations differ for weights given to individual models; see Wang et al., (2023) for an extensive review of the literature on forecast combinations.
@@
Although the equally weighted average does not consider correlations between forecast errors or the historical performance of individual forecasts, the literature suggests that it performs remarkably well compared to more complex combination techniques, becoming the most widely used combination rule (Bunn,, 1985; Genre et al.,, 2013).
Despite the effectiveness of the simple average, it is reasonable to assign a greater weight to the most accurate forecasting methods. Bates and Granger, (1969) introduced a method to determine the optimal weights by minimising the variance of the combined forecast error, assigning weights inversely proportional to the forecast error variance, given by
$\mathbf{w}_{t}^{\text{BG}}=\frac{\boldsymbol{\Sigma}_{t}^{-1}\mathbf{1}}{\mathbf{1}^{\prime}\boldsymbol{\Sigma}_{t}^{-1}\mathbf{1}},$ (10)
where $\boldsymbol{\Sigma}_{T+\tau|T}$ is the covariance matrix $N\times N$ of the $t$-step forecast errors.
@@ 4 Results
Table 1 reports summary statistics of carbon emissions by Italian electricity market zone for training, calibration, and test sets. At the aggregate level, average daily emissions in Italy range from 12,924.16 tCO2 (test set: January–December 2023) to 14,985.72 tCO2 (calibration set: July–December 2022). In line with Figure 1, a slight decrease in average emissions can be observed during 2023 compared to previous periods. At the zonal level, the North zone consistently records the highest mean emissions across all periods, ranging from 6,114.24 tCO2 (test set: January–December 2023) to 6,379.07 tCO2 (calibration set: July–December 2022), thus contributing substantially to national totals. Sicily and Sardinia show moderate average emission levels, characterised by relatively low variability, while the Centre-South and the South zones report higher average emissions than Sicily, but consistently lower than those of the North. In terms of variability (Stdev) there are no notable differences in the three periods.

### v2

@@ 3 Methodology
This section describes the setup of single models, forecast combinations, and the overall design of the forecast experiment. In the following, the hourly observations are denoted by $E_{1},\ldots,E_{t},\ldots,E_{n}$, where $t$ indexes the hours. To indicate observation at hour $h$ of day $i$, also the notation $E_{i,h}$ is used. The relationship between $t$, $i$, and $h$ is defined by t=24(i1)+ht=24(i-1)+h, with $t=1,\ldots,8760$, $i=1,\ldots,365$, and $h=1,\ldots,24$.
@@ 3.1 Single model setup
[…]
where: $B$ is the backshift operator, i.e. such that $B^{k}E_{t}=E_{t-k}$; $S$ denotes the seasonal period; $d$ and $D$ represent the orders of non-seasonal and seasonal differencing, respectively; is the vector of non-seasonal autoregressive coefficients; k is the vector of seasonal autoregressive coefficients; is a constant term; j is the vector of non-seasonal moving average coefficients; and l is the vector of seasonal moving average coefficients; 1, 2 and 3 represent the effects of day of the week $(W_{t})$, intra-annual seasonality $(Y_{t})$, and bank holidays $(B_{t})$, respectively; t is a white noise error term. SARIMAX models are identified by minimising their hourly average RMSE in the calibration set.
From a physical perspective, the processes involved in the production of carbon emissions from electricity generation are continuous in nature. Therefore, continuous-time stochastic processes offer a natural and potentially more accurate framework for modelling the evolution of carbon emissions. The adoption of functional models is motivated by the continuous nature of the underlying physical phenomenon (ramsay2006). The functional approach enables the recovery of information on emissions released within the intervals between discrete hourly observations, capturing intra-hour dynamics that would otherwise be overlooked. Carbon emissions from electricity generation can be naturally represented as curves, making it logical to capture the serial dependence among daily curves and their temporal dynamics (Damon2002; Chen2021; Damon2024). Figure 5 shows the daily Italian carbon emission curves for the period 2021-2023. The raw data were transformed into functional data using the 12 Fourier basis (ramsay2006).
[…]
@@
Although the SFARX model accounts for the serial correlation between the functional emission curves by modelling it as a linear operator in the functional space (bosq2000linear), non-parametric approaches such as GAM are also used to capture potential non-linearities and provide greater flexibility in modelling relationships (Hastie1986; hastie2015package). The GAM is defined by
[…]
The TBATS model, introduced by DeLivera2011, is another popular and flexible time series forecasting method, designed to handle multiple seasonalities, trend components, and short-term dependencies (Tawiah2023; stanislaus2024). The TBATS model (Hyndman2008; Hyndman2025) can be mathematically represented as follows:
[…]
Alongside TBATS, a "semi-functional" approach such as K-nearest neighbours (KNN) is also considered for its flexibility in leveraging similarity across daily carbon emissions profiles (Lora2007; Tajmouati2024). Given the carbon emissions recorded up to day i1i-1, the goal is to predict the daily emission profile for day $i$. Let $\mathbf{E}_{i}=(E_{i,1},\ldots,E_{i,h},\ldots,E_{i,24})$ denote the vector of hourly carbon emissions for a generic day $i$, where $E_{i,h}$ is the emission recorded at hour $h$ of day $i$. Let $D(i,j)$ define the discrepancy between the daily emission profiles of days $i$ and $j$. $D(i,j)$ is evaluated using the Euclidean distance as follows:
@@
The KNN method identifies the $k$ nearest neighbours of day i1i-1, forming the neighbour set $NS=\{\mathbf{E}^{(1)},\mathbf{E}^{(2)},\ldots,\mathbf{E}^{(k)}\}$, where $\mathbf{E}^{(1)}$ is the nearest day in the past, $\mathbf{E}^{(2)}$ the second nearest, and $\mathbf{E}^{(k)}$ the k-th nearest. The 24-hour emission profile for day $i$ is then predicted as a weighted average of the emissions on the days following those of the neighbour set:
@@
where $w_{k}\in[0,1]$ is a weight inversely proportional to the distance between $\mathbf{E}^{(k)}$ and $\mathbf{E}_{t-1}$. The optimal $k$ is chosen by minimising the hourly average RMSE (HNIN2024). An example is provided by the elbow plot in Figure 6 for Italy.
@@ 3.2 Forecast combination setup
The idea of combining forecasts dates back to BG1969, who demonstrated that forecast combinations typically outperform individual models (CLEMEN1989; surowiecki2005; TIMMERMANN2006). Basically, different forecast combinations differ for weights given to individual models; see WANG2023 for an extensive review of the literature on forecast combinations.
@@
Although the equally weighted average does not consider correlations between forecast errors or the historical performance of individual forecasts, the literature suggests that it performs remarkably well compared to more complex combination techniques, becoming the most widely used combination rule (BUNN1985151; GENRE2013108).
Despite the effectiveness of the simple average, it is reasonable to assign a greater weight to the most accurate forecasting methods. BG1969 introduced a method to determine the optimal weights by minimising the variance of the combined forecast error, assigning weights inversely proportional to the variance of the forecast error, given by
[…]
@@ 4 Results
Table B1 in the Appendix B reports summary statistics of carbon emissions by Italian electricity market zone for training, calibration, and test sets. At the aggregate level, average daily emissions in Italy range from 12,924.16 tCO2 (test set: January–December 2023) to 14,985.72 tCO2 (calibration set: July–December 2022). In line with Figure 2, a slight decrease in average emissions can be observed during 2023 compared to previous periods. At the zonal level, the North zone consistently records the highest mean emissions across all periods, ranging from 6,114.24 tCO2 (test set: January–December 2023) to 6,379.07 tCO2 (calibration set: July–December 2022), thus contributing substantially to national totals. Sicily and Sardinia show moderate average emission levels, characterised by relatively low variability, while the Centre-South and the South zones report higher average emissions than Sicily, but consistently lower than those of the North. In terms of variability (Stdev) there are no notable differences in the three periods.
@@
$\text{RMSE}=\frac{1}{24}\sum_{h=1}^{24}\text{RMSE}_{h}.$ (12)
$R^{2}=\frac{1}{24}\sum_{h=1}^{24}R^{2}_{\text{h}}.$ (14)
The DM test is used to assess the statistical significance of differences in forecast accuracy between individual models and forecast combinations (Diebold1995). This test provides a formal framework to compare the predictive performance of two competing models based on their forecast error differentials. The hypothesis system of the one-sided DM test is defined as follows:
[…]
The DM test statistic is calculated as
$DM=\frac{\bar{d}}{\sqrt{(1/T)\sum_{k=-m}^{m}{}_{k}}}$ (15)
where: $\bar{d}$ is the sample mean of the loss differential $d_{t,ij}$, defined as $d_{t,ij}=(e_{i,t}-e_{j,t})^{2}$ with $e_{i,t}$ and $e_{j,t}$ denoting the forecast errors of the models $i$ and $j$ at time $t$. The loss function $L(\cdot)$ is the squared error; k denotes the $k$-th order autocovariance of the loss differential series $d_{t,ij}$; $T$ is the number of forecast periods; $m$ is the maximum lag order used to estimate the long-run variance of $d_{t,ij}$. Under the null hypothesis $H_{0}$, the DM statistic asymptotically follows a standard normal distribution: $DM\sim\mathcal{N}(0,1)$.
[…]
$\begin{cases}H_{0}:d_{t,ij}=0\quad\forall i,j\\ H_{1}:d_{t,ij}\neq 0\text{ for some }i,j.\end{cases}$
The test statistic is defined as in equation (15) and for a current set of models $M$ can be aggregated into $DM_{M}=\max_{i,j}|DM_{ij}|$. Since the asymptotic distribution of $DM_{M}$ is nonstandard, the p-value is obtained using a bootstrap procedure (Hansen2011; bernardi2014).
4.1 Single model results
In the calibration test, the best performing models are selected by minimising the hourly average RMSE. The selected models are reported in Table 1.
For Italy, the selected models are SARIMAX(2,1,0)(1,1,1)24, FARX(2), GAM(3,1), and KNN(10). Table B3 in Appendix B reports the hourly RMSE, hourly average RMSE and R2, while the panel (a) in Figure 9 displays the relative RMSE compared to the Naïve model. The selected GAMs for Sicily and Sardinia are GAM(2,0) and GAM(1,1), respectively. These specifications are more parsimonious compared to GAM(3,1): GAM(2,0) excludes the weekly stochastic component and incorporates autoregressive effects up to 2 days, while GAM(1,1) includes the weekly stochastic component but limits autoregressive effects to 1 day.
Zone SARIMAX SFARX GAM KNN
Italy SARIMAX(2,1,0)(1,1,1)24 FARX(2) GAM(3,1) KNN(10)
North SARIMAX(3,1,0)(1,1,1)24 FAR(2) GAM(3,1) KNN(10)
Centre North SARIMAX(1,1,0)(1,1,0)24 FARX(2) GAM(3,1) KNN(35)
Centre South SARIMAX(3,1,0)(1,1,1)24 FARX(1) GAM(3,1) KNN(10)
South SARIMAX(1,1,0)(1,1,1)24 FARX(3) GAM(3,1) KNN(10)
Calabria SARIMAX(3,1,0)(1,1,1)24 FAR(2) GAM(3,1) KNN(15)
Sicily SARIMAX(2,1,0)(1,1,1)24 FARX(3) GAM(2,0) KNN(5)
Sardinia SARIMAX(1,1,0)(1,1,1)24 FARX(1) GAM(1,1) KNN(10)
For the test set, Table B2 in Appendix A reports the forecast performance of the best individual model of each modelling approach for Italy. RMSE values are provided for each hour of the day (0 to 23), along with the hourly average RMSE and the hourly average out of sample $R^{2}$. The GAM(3,1) model produces the lowest RMSE for most hours, especially during the daytime hours from 6:00 to 23:00. In contrast, the FARX(2) model performs best in the early morning hours from 0:00 to 5:00. The Naïve and TBATS models generally exhibit higher RMSE values. SARIMAX(2,1,0)(1,1,1)24 and KNN(10) models show moderate performance and remain behind the FARX(2) and GAM(3,1) models.
Figure 8 presents the residual diagnostics for the SARIMAX(2,1,0)(1,1,1)24 model estimated over the whole 2021-2023 period for Italy. The ACF and PACF (left and central panels) indicate that autocorrelations and partial autocorrelations remain within $\pm 0.07$ over the first 48 lags. The panel on the right displays the rolling mean and standard deviation, computed with one week overlapping window, showing no evidence of structural breaks over the sample. Furthermore, a Lagrange Multiplier test is performed to assess the presence of residual heteroscedasticity. The test yields a statistic of 20.552 with a p-value of 0.1137. The null hypothesis of no autoregressive conditional heteroscedasticity effects cannot be rejected. Overall, the residual analysis suggests that the weak stationarity assumption is satisfied, with residuals that exhibit no significant autocorrelation and approximately constant variance. The other market zones exhibit similar diagnostic results.
Figure 8: Residual diagnostics for the SARIMAX(2,1,0)(1,1,1)24 model estimated over the whole 2021-2023 period for Italy. The left panel displays the ACF of the residuals, the central panel reports the PACF, and the right panel shows the rolling mean and standard deviation computed with one week overlap window.
The hourly average RMSE and out-of-sample $R^{2}$ further confirm that the GAM(3,1) model is the best, followed by the FARX(2) and KNN(10) models. Although the GAM(3,1) model provides the most accurate and reliable forecasts during daytime hours, the FARX(2) model may be preferable in the early morning. These findings are further supported by the panel (b) in Figure 9, which displays the relative RMSE compared to the Naïve model for Italy in the test set.
Figures A1-A2 in Appendix A show the hourly performance of the best models across the seven market zones. Overall, GAM(3,1) consistently delivers the best performance in the North, Centre-North, Centre-South, South, and Calabria zones. In Sicily, FARX(3) is the best model, while in Sardinia, FARX(1) performs best, with only minor differences among the other models.
4.2 Forecast combination results
Table 2: Forecast combinations.
Model Method Combination
COMB1
SARIMAX, SFARX, GAM, TBATS, KNN
SA COMB1.SA
BG COMB1.BG
COMB2
SFARX, GAM
SA COMB2.SA
BG COMB2.BG
SEL COMB2.SEL
COMB3
SARIMAX, GAM
SA COMB3.SA
BG COMB3.BG
SEL COMB3.SEL
COMB4
GAM, KNN
SA COMB4.SA
BG COMB4.BG
SEL COMB4.SEL
Each combination combines the best performing models among those listed. BG, SA and SEL denote the combination method: BG refers to BG1969, SA denotes the simple average, and SEL indicates the selection-based combination method. For example, COMB1.BG for Italy combines the forecasts of SARIMAX(2,1,0)(1,1,1)24, FARX(2), GAM(3,1), TBATS and KNN(10), using BG1969. In contrast, COMB2.SEL combines the FARX(2) and GAM(3,1) forecasts using the selection-based approach.
The results for Italy are listed in Table B4 in the Appendix B, which reports the hourly average RMSE of the forecast combinations in the test set. The selection-based combination method produces lower hourly average RMSE and higher out-of-sample R2 compared to the simple average and BG1969. However, simple average combination yields a lower RMSE than the BG1969 for combinations including SFARX and GAM (COMB2.SA), as well as for the combination involving all models (COMB1.SA).
The best-performing method is COMB2.SEL, which selects the forecasts of the FARX(2) model during the early hours of the day (00:00-05:00) and those of the GAM(3,1) model for the remaining hours (06:00-23:00). This combination reduces RMSE by 45.26%, 15.09%, and 19.02% relative to the Naïve, GAM(3,1), and FARX(2) models, respectively.
The second-best combination, COMB3.SEL, achieves an hourly average RMSE of 1581.96, corresponding to a 35.21% reduction compared to the Naïve benchmark. Figure 10 shows for Italy the relative RMSE of the two top-performing forecast combinations, compared to the Naïve model.
Table 3 provides the DM test p-values, computed for the Italian CO2 prediction, comparing the forecast accuracy of the individual models with that of the two top-performing forecast combinations based on hourly average RMSE. Each p-value compares the model in the column ($j$) with the model in the column ($i$). The null hypothesis assumes equal forecast accuracy, whereas the alternative states that the column model $j$ yields more accurate forecasts than the row model $i$. In terms of individual models, the TBATS specification is significantly less accurate than all other single models and forecast combinations. Similarly, the SARIMAX(2,1,0)(1,1,1)24 model is significantly less accurate than all competitors except the TBATS model. The FARX(2) and KNN(10) models perform better than the SARIMAX(2,1,0)(1,1,1)24 and TBATS models. By contrast, FARX(2) and KNN(10) have forecast performances that are almost statistically equivalent. The GAM(3,1) model outperforms all other individual models, although it remains less accurate than all forecast combinations. COMB2.SEL provides the most accurate forecast combination, followed by COMB4.SEL. These results are consistent with the RMSE evaluation, confirming that forecast combinations improve accuracy beyond that of the best single models.
SARIMAXa FARX(2) GAM(3,1) TBATS KNN(10) COMB2.SEL COMB3.SEL
SARIMAXa - <0.05<0.05 <0.05<0.05 1.000 <0.05<0.05 <0.05<0.05 <0.05<0.05
FARX(2) 1.000 - <0.05<0.05 1.000 0.04 <0.05<0.05 <0.05<0.05
GAM(3,1) 1.000 1.000 - 1.000 1.000 <0.05<0.05 <0.05<0.05
TBATS <0.05<0.05 <0.05<0.05 <0.05<0.05 - <0.05<0.05 <0.05<0.05 <0.05<0.05
KNN10 1.000 0.96 <0.05<0.05 1.000 - <0.05<0.05 <0.05<0.05
COMB2.SEL 1.000 1.000 1.000 1.000 1.000 - 1.000
COMB3.SEL 1.000 1.000 1.000 1.000 1.000 <0.05<0.05 -
  • Note. a the model order is (2,1,0)(1,1,1)24.
Figure 10: Italy - Relative RMSE of the two top-performing forecast combinations (based on hourly average RMSE) compared to the Naïve model.
For the North zone, findings are in line with those obtained for Italy, as the North zone accounts for the largest share of electricity generation and associated carbon emissions. The optimal combination is COMB2.SEL, with an hourly average RMSE of 930.16, which is 46.88% lower than that of the Naïve model model. The second-best alternative has an RMSE that is 1.85% higher than that of COMB2.SEL.
The simple average can be considered the best combination method for the other market zones producing minimal or essentially equivalent errors compared to the others combinations methods (with numerically almost identical values). For the Centre-North zone, the optimal combination merges the functional autoregressive model with GAM, yielding an RMSE 22.93% lower than that of the Naïve model. By contrast, for the Centre-South, South, Calabria and Sicily, the best-performing combination pairs the SARIMAX model with GAM. These forecast combinations provide RMSE reductions of 28.54%, 24.57%, 25.32% and 20.55% relative to the Naïve model for the Centre-South, South, Calabria and Sicily, respectively.
For Sardinia, as expected, the best combination includes all models, since the individual models already perform well (Figure A2).
The DM test is also performed for the market zones, comparing the individual models with the best two forecast combination. The p-values of the DM test are reported in Tables B5B11 in Appendix B. The results confirm that the North zone aligns with the overall Italian outcomes, where the selection-based combination method performs best. In the other market zones, the simple average is the most effective combination method, producing minimal or equivalent forecast performance with the others methods.
[…]
For Italy and the North, the superior set includes only the forecast combination of the functional autoregressive model and the GAM. For the Centre-North, Centre-South, South, Calabria, and Sicily, the superior set includes the two best forecast combinations, which produce similar results. In Sardinia, the superior set includes the two best forecast combinations together with three individual models: FARX(1), TBATS, and SARIMAX(2,1,0)(1,1,1)24.
The last row of Table 4 reports the number of models excluded from the superior set for each market zone. Italy and the North exclude seven models, reflecting a more selective optimal set, while Centre-North, Centre-South, South, Calabria, and Sicily exclude six models each. In Sardinia, only three models are excluded, indicating that more models perform similarly well in this zone.
The MCS findings are consistent with those of the DM test. The simple average and the selection-based method perform best. Forecast combinations improve the accuracy of individual models. The most effective combinations include the functional autoregressive model and the GAM (Italy, North, Centre-North, and Calabria) or the SARIMAX model and the GAM (Centre-South, South, and Sicily). In Sardinia, where most individual models perform well, the best combination includes all models.
[…]
Rank Italy North Centre-North Centre-South South Calabria Sicily Sardinia
1
COMB2.SEL
COMB2.SEL
COMB2.SA
COMB3.SA
COMB3.SA
COMB2.SEL
COMB3.SA
COMB1.SA
2 - -
COMB2.SEL
COMB2.SEL
COMB2.SEL
COMB3.SA
COMB2.BG
COMB1.BG
3 - - - - - - -
FARX(1)
4 - - - - - - -
TBATS
5 - - - - - - -
SARIMAXa
Excl.
7 7 6 6 6 6 6 3
  • Note. a the model order is (2,1,0)(1,1,1)24. The last row reports the number of exuded models from the superior set of models.
5 Policy implications on electricity market scheduling
The results presented in the previous sections highlight the importance of selecting an appropriate method for predicting short-term CO2 emissions. Forecast performance varies across the electricity market zones, reflecting variations in local energy mixes. For instance, the outcomes for Italy and its mainland regions differ from those observed for the islands of Sicily and Sardinia. These differences are largely due to variations in energy sources and the degree of interconnection with the mainland, both of which affect electricity generation and associated CO2 emissions (Sapio2016; Sapio2020; Weinhold2021). Furthermore, the use of forecast combinations, which leverage the strengths of multiple individual models, has been shown to enhance predictive accuracy in most cases, particularly when the models capture different structural characteristics of the data (WANG2023)
These findings highlight the importance of considering zonal specificities when implementing flexible energy demand strategies (Abolghasemi2025). Short-term CO2 emissions forecasting enables flexible electricity demand scheduling to minimise emissions (BOKDE2021). For example, Figure 11 shows true ex-post (black) and predicted (red) carbon emissions using COMB2.SEL for Italy over 48 consecutive hours (1-3 February 2023). Mean emissions during this period are 16,345.42 tCO2 and 15,712.65 tCO2 for predicted and true values, respectively. The series starts at 12:00, when the day-ahead market is cleared; bids submitted before this time cover consumption for the following day (grey area between dashed blue lines). Flexible consumption is scheduled during hours with the lowest predicted emissions: 01:00-05:00 and 12:00-13:00. The red and black dots indicate the predicted and true minimum values, respectively.
Similar patterns have been reported for France, Germany, Norway, Denmark, and Poland (BOKDE2021), although the timing of flexible demand varies across countries. In Germany and France, demand is usually scheduled between 09:00 and 13:00, while in Denmark it occurs between 16:00 and 21:00. These differences underline the need to plan flexible demand according to each country and market area. Figure 12 shows the scheduling of 10 hours of flexible electricity demand in the Centre-South, illustrating that the timing of flexible energy demand windows can differ from those observed for Italy in Figure 11.
Figure 11: Example of scheduling 7 hours of flexible electricity consumption one day in advance in Italy. The red and black curves represent the fitted and true ex-post carbon emissions, respectively. The grey area indicates the day-ahead period. The minimum hours of the forecasted and true values are marked with red and black dots.
Figure 12: Example of scheduling 10 hours of flexible electricity consumption one day in advance in the Centre-South zone. The red and black curves represent the fitted and true ex-post carbon emissions, respectively. The grey area indicates the day-ahead period. The minimum hours of the forecasted and true values are marked with red and black dots.
6 Conclusions
This study investigates the short-term forecasting of carbon emissions from electricity generation in the Italian power market. Using hourly data from 2021 to 2023, several statistical models and forecast combination methods are evaluated and compared at the national and zonal levels. Four main model classes are considered: linear parametric models, functional parametric models, non-parametric and possibly non-linear models, and a semi-functional approach. Forecast combinations include the simple averaging, BG1969, and the selection-based approach.
For mainland Italy, including the national level, North, Centre-North, Centre-South, South, and Calabria, GAM provides the most accurate forecasts during daytime hours, whereas functional parametric models perform best during the early morning period. In general, GAM emerges as the most effective individual model based on the hourly average RMSE and DM test results. Among the forecast combination approaches, the selection-based method consistently performs best in Italy and the North, while the simple average is the most effective in the remaining market zones.
In Italy, the North, and the Centre-North, the most effective forecast combinations are those that employ the functional autoregressive model together with GAM. In contrast, in the Centre-South, South, Calabria, and Sicily, the best-performing combinations pair the SARIMAX model with GAM.
In Sardinia, all individual models deliver reasonably strong forecasting performance (Figure A2). As expected, the best-performing combination is the one that incorporates all models.
This work can be expanded in several ways. First, future work can integrate other approaches and models, such as neural networks and deep learning methods (Lowry2018; BOKDE2021; Maji2022b). Second, multi-day forecasts can also be explored, which can potentially allow for more advanced demand-side management strategies (Maji2022a; MASON2018; GIROLIMETTO2024a). Third, it is worth investigating the forecast reconciliation methods (Wickramasuriya2019; DIFONZO2024; girolimetto2024b), as they provide a coherent framework to ensure consistency between forecasts at different levels of aggregation. This is particularly important in the context of electricity-related CO2 emissions, where forecasts may be required at the zonal and national levels.
Declaration of interests
The authors report that financial support was provided by European Union - NextGenerationEU, Mission 4, Component 2, in the framework of the GRINS - Growing Resilient, INclusive and Sustainable project (GRINS PE00000018 - CUP C93C22005270001).
Acknowledgments
Data availability statement
Data are available from the authors.
Appendix A Figures
Appendix B Tables